OpenStudy (anonymous):

Find x and y so that the following equation is true \[\frac{2x}{3}+\frac{5yi}{4}=-12-25i\]

OpenStudy (anonymous):

Isolate the real and imaginary parts. \[\frac{2x}{3} = -12\]\[\frac{5yi}{4} = -25i\]

OpenStudy (anonymous):

x=-18 and y=-20?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

yes. x=-18 and y=20.

OpenStudy (anonymous):

x=-18 and y=-20

OpenStudy (foolaroundmath):

For x=-18, y = -20 to be true, it must be specified that x and y are real. Otherwise there are infinite number of solutions. Take for example, x = -3 and y = -20+4i